Devon Corporation uses a discount rate of 8% in its capital budgeting. Partial analysis of an investment in automated equipment with a useful life of 8 years has thus far yielded a net present value of −$502,541. This analysis did not include any estimates of the intangible benefits of automating this process nor did it include any estimate of the salvage value of the equipment. (Ignore income taxes.)

Required:

a. Ignoring any salvage value, how large would the additional cash flow per year from the intangible benefits have to be to make the investment in the automated equipment financially attractive? (Minimum Annual cash flows)

b. Ignoring any cash flows from intangible benefits, how large would the salvage value of the automated equipment have to be to make the investment in the automated equipment financially attractive? (Minimum Salvage Value)
Answer #1

A.additional cash flow per year required from intangible benefits = net present value required/Present value annuity factor

=502,541/PVAF(8%,8 Years)

= 502,541/5.7466

=$87,450.14

B. Required salvage value = desired NPV/present value factor

= 502,541/PVF(8,8 years)

= 502,541/0.5403

=$930,114.75